# Relaxation of first-class constraints and the quantization of gauge theories: from “matter without matter” to the reappearance of time in quantum gravity

### v1

@@ III.2.4 Deparametrization
A transformation of constraints that involves the solution of the original constraints in terms of a momentum variable is called “deparametrization.” [^19] This is conceptually the inverse procedure of parametrization, however, the two procedures may lead to theories that are not globally equivalent. See Footnote 18. The deparametrization in Eqs. (42) motivates a gauge fixing given by the conditions
$\bar{\chi}^{a}$ $\coloneqq\mathscr{Q}^{a}-f^{a}(\tau)\approx 0\ ,$ (46)
where $f^{a}$ are functions solely of $\tau$. These gauge conditions fix the first $P$ multipliers to be [cf. Eq. (25)]:
@@
$S_{\rm new}^{\rm par}$ $=\int_{\tau_{0}}^{\tau_{1}}\!\mathrm{d}\tau\left[\mathcal{A}_{\mu}(z)\dot{z}^{\mu}-H_{\rm red}\right]\,,$ (48)
@@
This corresponds to the action functional of a system with Hamiltonian $H_{\rm red}$, understood as a function solely of time and $z$ [^20] Recall that all constants of Nature that were promoted to fields have been replaced by the corresponding $-\mathscr{H}_{a}$, and their conjugate cyclic variables were replaced by $f^{a}(\tau)$. Of course, $H_{\rm red}$ may still depend on other constant parameters that are not written explicitly. Thus, the solution of the constraints and the gauge fixing in Eqs. (41) and (46) effectively yield a system with a reduced set of degrees of freedom [the $z$ fields instead of the set $(z,\mathscr{Q},\mathscr{P})$]. For this reason, the Hamiltonian $H_{\rm red}$ is sometimes referred to as the ‘reduced Hamiltonian’ Kiefer:book . Likewise, the gauge freedom related to the constraints $\phi_{a}$ ($a=1,\ldots,P$) and its associated “gauge parameters” that change with gauge transformations have been removed: that is why one says that the system has been (partially) “deparametrized,” and Eqs. (42) are the deparametrized form of the constraints $\phi_{a}$ ($a=1,\ldots,P$). If we can set $f^{a_{0}}(\tau)=\tau$, then the cyclic field $\mathscr{Q}^{a_{0}}$ plays the role of a physical clock. We will see in Sec. III.2.6 how this is related to the Fock–Stueckelberg approach, and we discuss such a deparametrization in the context of general relativity in Secs. III.7.4 and IV.
@@ III.2.6 Equivalence with the Fock–Stueckelberg theory
because $f^{a}(\tau)$ are functions solely of $\tau$, and $z$ Poisson-commutes with $\mathscr{P}$ by definition. The equations (50) are exactly those derived from the action in Eq. (48). In this way, as far as the $z$ fields are concerned, both the Fock–Stueckelberg and constants of Nature approaches are equivalent. This is of course expected because the deparametrization via constants of Nature induced a choice of gauge [cf. Eqs. (46) and (47)], which can then be used to seed the Fock–Stueckelberg approach. However, it is important to emphasize that this equivalence is established starting from the action in Eq. (45), in which the constraints are already in deparametrized form. As this form holds only locally in general, the Fock–Stueckelberg theory defined from a more general parent theory, with action given in Eq. (40) for example, may generally exhibit global differences. Whereas the physical meaning of the constants of Nature approach becomes most clear with deparametrization, [^21] In the constants of Nature approach, the effects of the relaxation of constraints, as well as the resulting physical equations of motion derived from the action in Eq. (48), are most clearly seen if the deparametrization can be explicitly performed. the effects of the Fock–Stueckelberg approach on the relaxation of constraints are evident without deparametrization, as they only require a well-defined noncanonical gauge condition.
@@ III.3 First-class subalgebras
Let us consider a parent gauge theory with the following property: we can divide the constraints into $\phi=(u,K)$ such that $u_{a}=\phi_{a}$ ($a=1,\ldots,N_{u}$) and $K_{a}=\phi_{N_{u}+a}$ ($a=1,\ldots,N-N_{u}$), and we assume that the structure functions have the property $\mathcal{V}^{c}_{b}=0,\mathcal{F}_{ab}^{c}=0$ if $c=1,\ldots,N_{u}$. If this is not the case, it may be possible to bring the structure functions into this form by a redefinition of the constraints (cf. Sec. III.2.3), [^22] In local regions of field space, it may be possible to choose $\tensor{M}{{}_{a}^{b}}$ in $\bar{\phi}_{a}=\tensor{M}{{}_{a}^{b}}\phi_{b}$ (cf. Sec. III.2.3) so that the new structure functions vanish, and the new constraints locally satisfy an Abelian algebra $\{\bar{\phi}_{a},\bar{\phi}_{b}\}=0$. In general, however, it is not possible to achieve this ‘local Abelianization’ in a global manner HT:book . and the $u$ constraints Poisson-commute if the $K$ constraints are satisfied: [^23] Indices contracted with the index of $\phi_{a}$ are summed from $1$ to $N$. However, indices contracted with the index of $u_{a}$ are to be summed from $1$ to $N_{u}$ only, whereas indices contracted with the index of $K_{a}$ are to be summed from $1$ to $N-N_{u}$.
@@
on the hypersurface defined by $\chi^{a}\approx 0$. If we also enforce $K_{b}\approx 0$, then Eq. (54) can be solved for $\lambda^{a}$ ($a=1,\ldots,N_{u}$) similarly to Eq. (25) by fixing the first $N_{u}$ multipliers to be solely functions of $\tau$:
$\lambda^{a}=\dot{f}^{a}(\tau)\ ,(a=1,\ldots,N_{u})\ .$ (55)
@@
The corresponding Fock–Stueckelberg theory with the noncanonical gauge conditions in Eq. (55) leads to the constraints $\lambda^{a}\approx\dot{f}^{a}(\tau),u_{a}+w_{a}\approx 0,K_{b}\approx 0$ [cf. Eqs. (29)]. Due to the conservation of $u_{a}$ on the hypersurface defined by $K_{b}\approx 0$, the effect of the Fock–Stueckelberg constraint $u_{a}+w_{a}\approx 0$ is simply to assign a constant value to the $w_{a}$ multiplier, and the constraints $u_{a}$ are relaxed: instead of being forced to take the value 00, they can now take any constant value. This is useful because many models of interest, such as (bosonic) relativistic particles (which have $H_{0}\equiv 0$ and an Abelian algebra of constraints), can be seen as particular cases of this example.
@@ III.5 To gauge or not to gauge time?
$S_{\rm mech}\coloneqq\int_{\tau_{0}}^{\tau_{1}}\mathrm{d}\tau\ \left[\mathcal{A}_{\mu}(z)\dot{z}^{\mu}-H(z)\right]\ ,$ (61)
which forms the starting point of the usual Hamilton variational principle. The particle’s degrees of freedom, given by the $z^{\mu}$ ($\mu=1,\ldots,Z$) variables, are scalar fields on its worldline, and $\tau$ is an ‘absolute’ or ‘external’ time parameter that cannot be defined from $z$. Since the Hamiltonian is assumed to have no explicit time dependence, it is conserved, and time translations form a global (rigid) symmetry. They can be ‘gauged’ (i.e., turned into a local symmetry) if we introduce a worldline scalar density $\lambda(\tau)$ as an independent auxiliary field, $H(z)\to\lambda H(z)$, so as to make both terms in the Lagrangian (the symplectic potential term and the Hamiltonian) transform as worldline scalar densities. The new action
$\tilde{S}_{\rm mech}\coloneqq\int_{\tau_{0}}^{\tau_{1}}\mathrm{d}\tau\ \left[\mathcal{A}_{\mu}(z)\dot{z}^{\mu}-\lambda H(z)\right]$ (62)
@@
$S_{\rm FS}\coloneqq\int_{\tau_{0}}^{\tau_{1}}\mathrm{d}\tau\ \left[\mathcal{A}_{\mu}(z)\dot{z}^{\mu}-\lambda H(z)-w(\lambda-\Lambda)\right]\ ,$ (63)
@@
$\tilde{S}_{\rm FS}\coloneqq\int_{\tau_{0}}^{\tau_{1}}\mathrm{d}\tau\ \left[\mathcal{A}_{\mu}(z)\dot{z}^{\mu}-\Lambda H(z)\right]\ .$ (64)
@@ III.5.1 The relativistic particle: from Jacobi to Hamilton and back
where $\sigma=\pm 1$ is the sign of $\lambda$, which plays the role of the ‘einbein’ on the particle’s worldline. If the function $f_{a}(q)\equiv 0$, and with $\mathcal{E}$ being a constant, this is precisely the usual Jacobi action (up to an overall sign) that can be used to define time-repametrization invariant trajectories in classical mechanics via its extremization (the Jacobi action principle)Lanczos . [^28] Starting from the Arnowitt–Deser–Misner (ADM) formulation of general relativity [discussed in Kiefer:book ; ADM (and references therein) and briefly in Sec. III.7.2], a similar procedure leads to a square-root action for general relativity, known as the Baierlein–Sharp–Wheeler action Kiefer:book ; Baierlein . Analogously, starting from the Polyakov action for a relativistic string, one can obtain a square–root action known as the Nambu–Goto action strings1 ; strings2 ; strings3 ; strings4 . Furthermore, if $\sigma=1$, $f_{a}=-eA_{a}$ is constant, $V(q)\equiv 0$ and $\mathcal{E}\equiv-mc^{2}/2$ is constant, the action in Eq. (68) reduces to the usual action for a relativistic particle with charge $e$ coupled to a constant external electromagnetic field. In this case, the corresponding Hamilton–Jacobi equation reads
@@
It is also worthwhile to notice that the Fock–Stueckelberg trajectories of a relativistic particle (with $V\equiv 0$) need not be timelike or null throughout configuration space. Only those trajectories satisfying the initial-value Hamiltonian constraint will conform to such a causal restriction. In this way, the transition from the parent to the Fock–Stueckelberg theory allows more solutions to the equations of motion: whereas the parent theory constrains initial-values, these are free in the Fock–Stueckelberg theory. This was one of the main motivations of Stueckelberg, as the relaxation of this constraint meant that one could consider trajectories going forwards and backwards in Minkowski time, and this could be useful in the description of pair creation and annihiliation Stueck ; Stueck1 ; Stueck2 .
@@
This is an unconstrained action of the type of the usual Hamilton principle. Its global time-translation invariance leads exactly to the conservation law expressed in Eq. (73), which is what one obtains with the Fock–Stueckelberg theory. Thus, the Fock–Stueckelberg approach leads to a relaxation of first-class constraints by adding new degrees of freedom, which generally can be seen as the $w$ multipliers [e.g., as in Eq. (63)], but, in certain cases, these new variables can be identified with constants of motion (promoted to physical conserved fields), as it was discussed in Sec. III.2, and as we see here in the case of the relativistic particle, where $\mathcal{E}$ is related to the particle’s rest energy.
@@ III.5.3 The work of Fock and Stueckelberg
Similarly, Stueckelberg proposed in Stueck ; Stueck1 ; Stueck2 a “new relativistic mechanics” by positing a new action principle for the relativistic particle. The proposed action was a particular case of Eq. (70), and Stueckelberg stressed that the new formulation could be straightforwardly quantized à la Schrödinger, and it could also be used to describe the creation and annihilation of particles not only in the quantum theory but also in the classical theory [due to the fact that the trajectories that can be derived from Eq. (70) need not be always timelike or null if the particle is interacting with an external field]. [^34] The works Stueck ; Stueck1 ; Stueck2 of Stueckelberg have a conceptual similarity with but are distinct from the other famous Stueckelberg method StueckTrick ; StueckTrick1 : that of introducing a scalar field to preserve gauge invariance in a massive Abelian gauge theory of the Yang–Mills type. In particular, the particle’s mass was to be understood as an integration constant associated to each worldline (an idea which would later make a comeback concerning the cosmological constant in unimodular gravity, and possibly any constant of Nature in quantum gravity Magueijo1 ; Magueijo2 ; Magueijo3 ; Magueijo4 ; Magueijo5 ; Magueijo6 ; Magueijo7 , cf. Secs. III.2 and IV).
@@ III.6 Electrodynamics and “charge without charge”
Dirac’s electrodynamics inspired many further works (e.g., post ; post1 ; post11 ; post2 ; post3 ), [^39] It is worthwhile to mention that Nambu NambuEM verified that the corresponding quantum field theory and $S$-matrix of Dirac’s new electrodynamics with a timelike four-vector potential (with $k>0$) agree with the usual results of quantum electrodynamics at least up to the lowest orders in perturbation theory. More precisely, Nambu considered the Lagrangian $-\frac{1}{4}F^{ab}F_{ab}-j_{a}A^{a}$, where $j^{a}$ is a conserved current, and the gauge condition is tacitly already satisfied; i.e., $A_{0}$ is defined as $A_{0}\coloneqq\pm\sqrt{A_{1}^{2}+A_{2}^{2}+A_{3}^{2}+k^{2}}$. This corresponds to Dirac’s new electrodynamics in the presence of some four-current $j^{a}$ and with the gauge condition already imposed at the level of the action [similarly to Eq. (31)]. including an extension to non-Abelian gauge theories in post11 , and the idea was recently reincarnated in Kaplan2 , which is part of the series of works Kaplan ; Kaplan1 ; Kaplan2 , where the argument that the modified theory leads to a more straightforward quantization is again recorded. However, both Stueckelberg and Dirac make it clear that one is producing a different classical theory which is to be quantized, rather than arguing that the modified theory follows from a consistent quantization or concerns the “classical equations of motion of quantized gauge theories” (as in Kaplan1 ; Kaplan2 ). The Fock–Stueckelberg approach does not concern the consistency of quantization methods of gauge theories, [^40] One such method is the powerful BRST cohomology HT:book , which does not require that a gauge be fixed prior to quantization in order to achieve a nonambiguous Hamiltonian. but it rather offers an ingenious way to classically add new degrees of freedom by breaking the original gauge symmetry. [^41] Of course, Nature is presumably “quantum first,” and classical behavior is relegated to an approximate limit. Even so, this does not imply that, in order to be well defined, the fundamental quantum theory must break gauge symmetries, which would then only emerge in particular cases. We will further discuss this in Sec. V.

### v2

@@ III.2.4 Deparametrization
A transformation of constraints that involves the solution of the original constraints in terms of a momentum variable is called “deparametrization.” [^19] This is conceptually the inverse procedure of parametrization. However, the two procedures may lead to theories that are not globally equivalent. See Footnote 18. The deparametrization in Eqs. (42) motivates a gauge fixing given by the conditions
$\bar{\chi}^{a}$ $\coloneqq\mathscr{Q}^{a}-f^{a}\approx 0\ ,$ (46)
where $f^{a}$ are functions solely of the spacetime point. These gauge conditions fix the first $P$ multipliers to be [cf. Eq. (25)]:
@@
$S_{\rm new}^{\rm par}$ $=\int_{\tau_{0}}^{\tau_{1}}\!\mathrm{d}\tau\left[\mathcal{A}_{\mu}\dot{z}^{\mu}-H_{\rm red}\right]\,,$ (48)
@@
This corresponds to the action functional of a system with Hamiltonian $H_{\rm red}$, understood as a function solely of the spacetime point and $(z,\ldots,\partial^{n}z)$. [^20] Recall that all constants of Nature that were promoted to fields have been replaced by the corresponding $-\mathscr{H}_{a}$, and their conjugate cyclic variables were replaced by $f^{a}$. Of course, $H_{\rm red}$ may still depend on other constant parameters that are not written explicitly. Thus, the solution of the constraints and the gauge fixing in Eqs. (41) and (46) effectively yield a system with a reduced set of degrees of freedom [the $z$ fields instead of the set $(z,\mathscr{Q},\mathscr{P})$]. For this reason, the Hamiltonian $H_{\rm red}$ is sometimes referred to as the ‘reduced Hamiltonian’ Kiefer:book . Likewise, the gauge freedom related to the constraints $\phi_{a}$ ($a=1,\ldots,P$) and its associated “gauge parameters” that change with gauge transformations have been removed: that is why one says that the system has been (partially) “deparametrized,” and Eqs. (42) are the deparametrized form of the constraints $\phi_{a}$ ($a=1,\ldots,P$). If we can set $f^{a_{0}}=\tau$, then the field $\mathscr{Q}^{a_{0}}$ plays the role of a physical clock. We will see in Sec. III.2.6 how this is related to the Fock–Stueckelberg approach, and we discuss such a deparametrization in the context of general relativity in Secs. III.7.4 and IV.
@@ III.2.6 Equivalence with the Fock–Stueckelberg theory
because $f^{a}$ are functions solely of the spacetime point, and $z$ Poisson-commutes with $\mathscr{P}$ by definition. The equations (50) are exactly those derived from the action in Eq. (48). In this way, as far as the $z$ fields are concerned, both the Fock–Stueckelberg and constants of Nature approaches are equivalent. This is of course expected because the deparametrization via constants of Nature induced a choice of gauge [cf. Eqs. (46) and (47)], which can then be used to seed the Fock–Stueckelberg approach. However, it is important to emphasize that this equivalence is established starting from the action in Eq. (45), in which the constraints are already in deparametrized form. As this form holds only locally in general, the Fock–Stueckelberg theory defined from a more general parent theory, with action given in Eq. (40) for example, may generally exhibit global differences. Whereas the physical meaning of the constants of Nature approach becomes most clear with deparametrization, [^21] In the constants of Nature approach, the effects of the relaxation of constraints, as well as the resulting physical equations of motion derived from the action in Eq. (48), are most clearly seen if the deparametrization can be explicitly performed. the effects of the Fock–Stueckelberg approach on the relaxation of constraints are evident without deparametrization, as they only require a well-defined noncanonical gauge condition.
@@ III.3 First-class subalgebras
Let us consider a parent gauge theory with the following property: we can divide the constraints into $\phi=(u,K)$ such that $u_{a}=\phi_{a}$ ($a=1,\ldots,N_{u}$) and $K_{a}=\phi_{N_{u}+a}$ ($a=1,\ldots,N-N_{u}$), and we assume that the structure functions have the property $\mathcal{V}^{c}_{b}=0,\mathcal{F}_{ab}^{c}=0$ if $c=1,\ldots,N_{u}$. If this is not the case, it may be possible to bring the structure functions into this form by a redefinition of the constraints (cf. Sec. III.2.3), [^22] In local regions of field space, it may be possible to choose $\tensor{M}{{}_{a}^{b}}$ in $\bar{\phi}_{a}=\tensor{M}{{}_{a}^{b}}\phi_{b}$ (cf. Sec. III.2.3) so that the new structure functions vanish, and the new constraints locally satisfy an Abelian algebra $\{\bar{\phi}_{a},\bar{\phi}_{b}\}=0$. In general, however, it is not possible to achieve this ‘Abelianization’ in a global manner HT:book . and the $u$ constraints Poisson-commute if the $K$ constraints are satisfied: [^23] Indices contracted with the index of $\phi_{a}$ are summed from $1$ to $N$. However, indices contracted with the index of $u_{a}$ are to be summed from $1$ to $N_{u}$ only, whereas indices contracted with the index of $K_{a}$ are to be summed from $1$ to $N-N_{u}$.
@@
on the hypersurface defined by $\chi^{a}\approx 0$. If we also enforce $K_{b}\approx 0$, then Eq. (54) can be solved for $\lambda^{a}$ ($a=1,\ldots,N_{u}$) similarly to Eq. (25) by fixing the first $N_{u}$ multipliers to be:
$\lambda^{a}=\dot{f}^{a}(\tau,\mathbf{x})\ ,\ (a=1,\ldots,N_{u})\ .$ (55)
@@
The corresponding Fock–Stueckelberg theory with the noncanonical gauge conditions in Eq. (55) leads to the constraints $\lambda^{a}\approx\dot{f}^{a}(\tau,\mathbf{x}),u_{a}+w_{a}\approx 0,K_{b}\approx 0$ [cf. Eqs. (29)]. Due to the conservation of $u_{a}$ on the hypersurface defined by $K_{b}\approx 0$, the effect of the Fock–Stueckelberg constraint $u_{a}+w_{a}\approx 0$ is simply to assign a constant-in-time value to the $w_{a}$ multiplier, and the constraints $u_{a}$ are relaxed: instead of being forced to take the value 00, they can now take any constant-in-time value. This is useful because many models of interest, such as (bosonic) relativistic particles (which have $d=0,H_{0}\equiv 0$ and an Abelian algebra of constraints), can be seen as particular cases of this example.
@@ III.5 To gauge or not to gauge time?
$S_{\rm mech}\coloneqq\int_{\tau_{0}}^{\tau_{1}}\mathrm{d}\tau\ \left[\mathcal{A}_{\mu}\dot{z}^{\mu}-H(z)\right]\ ,$ (61)
which forms the starting point of the usual Hamilton variational principle. The particle’s degrees of freedom, given by the $z^{\mu}$ ($\mu=1,\ldots,Z$) variables, are scalar fields on its worldline, and $\tau$ is an ‘absolute’ or ‘external’ time parameter. Since the Hamiltonian is assumed to have no explicit time dependence, it is conserved, and time translations form a global (rigid) symmetry. They can be ‘gauged’ (i.e., turned into a local symmetry) if we introduce a worldline scalar density $\lambda(\tau)$ as an independent auxiliary field, $H(z)\to\lambda H(z)$, so as to make both terms in the Lagrangian (the symplectic potential term and the Hamiltonian) transform as worldline scalar densities. The new action
$\tilde{S}_{\rm mech}\coloneqq\int_{\tau_{0}}^{\tau_{1}}\mathrm{d}\tau\ \left[\mathcal{A}_{\mu}\dot{z}^{\mu}-\lambda H(z)\right]$ (62)
@@
$S_{\rm FS}\coloneqq\int_{\tau_{0}}^{\tau_{1}}\mathrm{d}\tau\ \left[\mathcal{A}_{\mu}\dot{z}^{\mu}-\lambda H(z)-w(\lambda-\Lambda)\right]\ ,$ (63)
@@
$\tilde{S}_{\rm FS}\coloneqq\int_{\tau_{0}}^{\tau_{1}}\mathrm{d}\tau\ \left[\mathcal{A}_{\mu}\dot{z}^{\mu}-\Lambda H(z)\right]\ .$ (64)
@@ III.5.1 The relativistic particle: from Jacobi to Hamilton and back
where $\sigma=\pm 1$ is the sign of $\lambda$, which plays the role of the ‘einbein’ on the particle’s worldline. If the function $f_{a}(q)\equiv 0$, and with $\mathcal{E}$ being a constant, this is precisely the usual Jacobi action (up to an overall sign) that can be used to define time-reparametrization invariant trajectories in classical mechanics via its extremization (the Jacobi action principle)Lanczos . [^28] Starting from the Arnowitt–Deser–Misner (ADM) formulation of general relativity [discussed in Kiefer:book ; ADM (and references therein) and briefly in Sec. III.7.2], a similar procedure leads to a square-root action for general relativity, known as the Baierlein–Sharp–Wheeler action Kiefer:book ; Baierlein . Analogously, starting from the Polyakov action for a relativistic string, one can obtain a square–root action known as the Nambu–Goto action strings1 ; strings2 ; strings3 ; strings4 . Furthermore, if $\sigma=1$, $f_{a}=-eA_{a}$, $V(q)\equiv 0$ and $\mathcal{E}\equiv-mc^{2}/2$, the action in Eq. (68) reduces to the usual action for a relativistic particle with charge $e$ coupled to an external electromagnetic field. In this case, the corresponding Hamilton–Jacobi equation reads
@@
It is also worthwhile to notice that the Fock–Stueckelberg trajectories of a relativistic particle (e.g., with $V\neq 0$) need not be timelike or null throughout configuration space. Indeed, the transition from the parent to the Fock–Stueckelberg theory allows more solutions to the equations of motion: whereas the parent theory constrains initial-values, these are free in the Fock–Stueckelberg theory. This allowed Stueckelberg to consider trajectories going forwards and backwards in Minkowski time, which was useful in the description of pair creation and annihilation Stueck ; Stueck1 ; Stueck2 .
@@
This is an unconstrained action of the type of the usual Hamilton principle. Its global time-translation invariance leads exactly to the conservation law expressed in Eq. (73), which is what one obtains with the Fock–Stueckelberg theory. Thus, the Fock–Stueckelberg approach leads to a relaxation of first-class constraints by adding new degrees of freedom, which generally can be seen as the $w$ multipliers [e.g., as in Eq. (63)], but, in certain cases, these new variables can be identified with constants of motion (promoted to physical conserved fields), as it was discussed in Sec. III.2, and as we see here in the case of the relativistic particle, where $\mathcal{E}$ may be related to the particle’s rest energy.
@@ III.5.3 The work of Fock and Stueckelberg
Similarly, Stueckelberg proposed in Stueck ; Stueck1 ; Stueck2 a “new relativistic mechanics” by positing a new action principle for the relativistic particle. The proposed action was a particular case of Eq. (70), and Stueckelberg stressed that the new formulation could be straightforwardly quantized à la Schrödinger, and it could also be used to describe the creation and annihilation of particles not only in the quantum theory but also in the classical theory [due to the fact that the trajectories that can be derived from Eq. (70) need not be always timelike or null]. [^34] The works Stueck ; Stueck1 ; Stueck2 of Stueckelberg have a conceptual similarity with but are distinct from the other famous Stueckelberg method StueckTrick ; StueckTrick1 ; StueckTrick2 : that of introducing a scalar field to preserve gauge invariance in a massive Abelian gauge theory of the Yang–Mills type. The value of the Hamiltonian (or, in the particular case of a free particle, the particle’s mass) was to be understood as an integration constant associated to each worldline (an idea which would later make a comeback concerning the cosmological constant in unimodular gravity, and possibly any constant of Nature in quantum gravity Magueijo1 ; Magueijo2 ; Magueijo3 ; Magueijo4 ; Magueijo5 ; Magueijo6 ; Magueijo7 , cf. Secs. III.2 and IV).
@@ III.6 Electrodynamics and “charge without charge”
Dirac’s electrodynamics inspired many further works (e.g., post ; post1 ; post11 ; post2 ; post3 ), [^39] It is worthwhile to mention that Nambu NambuEM verified that the corresponding quantum field theory and $S$-matrix of Dirac’s new electrodynamics with a timelike four-vector potential (with $k^{2}>0$) agree with the usual results of quantum electrodynamics at least up to the lowest orders in perturbation theory. More precisely, Nambu considered the Lagrangian $-\frac{1}{4}F^{ab}F_{ab}-j_{a}A^{a}$, where $j^{a}$ is a conserved current, and the gauge condition is tacitly already satisfied; i.e., $A_{0}$ is defined as $A_{0}\coloneqq\pm\sqrt{A_{1}^{2}+A_{2}^{2}+A_{3}^{2}+k^{2}}$. This corresponds to Dirac’s new electrodynamics in the presence of some four-current $j^{a}$ and with the gauge condition already imposed at the level of the action [similarly to Eq. (31)]. including an extension to non-Abelian gauge theories in post11 , and the idea was recently reincarnated in Kaplan2 , which is part of the series of works Kaplan ; Kaplan1 ; Kaplan2 , where the argument that the modified theory leads to a more straightforward quantization is again recorded. However, both Stueckelberg and Dirac make it clear that one is producing a different classical theory which is to be quantized, rather than arguing that the modified theory follows from a consistent quantization or concerns the “classical equations of motion of quantized gauge theories” (as in Kaplan1 ; Kaplan2 ). The Fock–Stueckelberg approach does not concern the consistency of quantization methods of gauge theories, [^40] One such method is the powerful BRST cohomology HT:book , which does not require that a gauge be fixed prior to quantization in order to achieve a nonambiguous Hamiltonian. but it rather offers an ingenious way to classically add new degrees of freedom by breaking the original gauge symmetry. [^41] Of course, Nature is presumably “quantum first,” and classical behavior is relegated to an approximate limit. Even so, this does not imply that, in order to be well defined, the fundamental quantum theory must break gauge symmetries, which would then only emerge in particular cases. We will further discuss this in Sec. V.